Please help me solve
Two angles are supplementary. One angle is 15 degrees more than twice the other. Find the measure of each angle.

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

Two angles are supplementary. One angle is 15 degrees more than twice the other. Find the measure of each angle.
-------------
A + B = 180
A = 2B + 15

Answer by rfer(16322)

x+2x+15=180
3x=180-15
3x=165
x=165/3
x=55
2x+15=125